A jet flying at 133 m/s banks to make a horizontal circular turn. The radius of the turn is 3810 m, and the mass of the jet is 1.84 105 kg. Calculate the magnitude of the necessary lifting force.
Answer #1

given
v = 133 m/s
r = 3810 m
m = 1.84*10^5 kg

centripetal force needed, F_centripetal = m*v^2/r

= 1.84*10^5*133^2/3810

= 854272 N

Force required to balance gravity, Fy = Fg

= m*g

= 1.84*10^5*9.8

= 1803200 N

Total magnitude of lifting force, F_lifting = sqrt(F_centripetal^2 + Fy^2)

= sqrt(854272^2 + 1803200^2)

= 2.00*10^6 N <<<<<<<-----------------Answer